Pixel:

what percentage is 110.72 out of 120

Pixel:

@dude

dude:

Divide the numbers and then multiply by 100

dude:

\(\large{\frac{110.72}{120}}\cdot 100\)

Pixel:

108%

dude:

What is \(\frac{110.72}{120}\) o.O

dude:

It isnt 108

Pixel:

92%

dude:

92.26% Yeh rounded would be 92